### v1

@@ Abstract
We present a pair-wise force law in a system of $N$ particles that produces analytic solutions for arbitrary number of particles, masses, and initial conditions. Each pair of particles interacts via a force that is proportional to the product of their masses and their separation distance, with the force directed radially. We show that, despite the $N$-body interaction, each particle behaves as if it interacts only with the center of mass of the system. This effective two-body interaction behaves as Hooke’s Law with a common frequency for all particles, with the familiar analytic solutions for the trajectories. With these analytic solutions, it is possible to efficiently simulate a collection of these particles and incorporate other external forces. As an example, we simulate the particles within an adiabatically expanding container and calculate pressure and temperature in both the attractive and repulsive cases.
1 Introduction
@@
The general $N$-body problem has long been a challenge in analytical mechanics. Newtonian gravity has been of particular interest, where the three-body problem admits analytical solutions only in special cases [1, 2, 3, 4, 5, 6, 7, 8, 9] and interactions between a sphere and a hoop generates intricate orbital motions [10]. Even with numerical solutions, the chaotic nature of $N$-body problems means that arbitrary systems will be unstable, presenting pedagogical challenges for students interested in simulating such a system.
This paper presents an investigation of a pairwise force law with analytic solutions in an $N$-body system. Intriguingly, this force law is shown to be equivalent to a central force acting on each particle, and each individual particle exhibits motion independent of its immediate neighbors.
@@
Given their mathematical simplicity, these solutions can be conveniently integrated into undergraduate courses on advanced mechanics, within units on harmonic oscillators or gravitational interactions. Moreover, due to their computational efficiency, these systems can be accurately simulated with large $N$ values on personal computers, thus rendering them suitable for inclusion in computational projects or courses on numerical methods.
@@
The absence of short range interactions, while convenient for quantum systems based on similar interactions [11, 12, 13, 14, 15], presents a lack of dynamism in the classical case. For this reason, further applications of this force law are most intriguing when paired with an external interaction. In this paper, we explore a thermodynamic application with the system in an expanding container, and suggest follow-up student projects.
@@
The paper is organized as follows. In Section 2 the force law is introduced, and equivalence of an $N$-body interaction to a 2-body interaction with the center of mass is shown. In Section 3, we present thermodynamic results for attractive and repulsive systems in an expanding container. Conclusions are provided in Section 4. Supplemental materials include a number of short python programs. These include one for plotting the orbits of an N-particle system given a specified set of mass and initial condition values based on the analytic trajectories; a similar program that runs a numerical simulation of a similar system directly from the pair-wise forces. The collision algorithm is presented in Appendix A, while ideas for possible research projects for advanced undergraduate students involving these two systems are discussed in Appendix B.
2 Force Law and General Solution
@@
Bertrand’s theorem shows that the only central forces that guarantee closed orbits are the inverse square and linear force laws [16]. Inspired by this, we introduce a toy model of a linear gravity-like force
@@
2.1 Solutions for point particles
@@
where $M=\sum_{k}m_{k}$ is the total mass of the system, and $\mathbf{R}_{\mathrm{cm}}$ is the position of the center of mass, which need not be stationary. Thus, the particle behaves as if it is governed by Hooke’s Law centered at $\mathbf{R}_{\mathrm{cm}}$, with effective spring constant $k_{i}=JMm_{i},$ as illustrated in Figure 1.
@@
where $\omega=\sqrt{JM}$ for all particles, and the constant vectors $\mathbf{A},\,\mathbf{B}$ are determined as usual from initial conditions. In the absence of external forces, the center of mass of the system moves at constant velocity, and we can use coordinates with $\mathbf{R}_{\mathrm{cm}}=0$ so that the particles move in closed elliptical orbits with a common frequency, consistent with Bertrand’s Theorem [3, 17, 18, 19].
Thus, the $N$-body interaction has an analytical solution, with the motion of each particle effectively described as an interaction with the center of mass of the system, and the trajectories are described by analytical expressions. This differs from the general $N$-body problem, which cannot be solved analytically. Additionally, particles in this system remain bound, in contrast to the inverse-square law, which can eject particles from the system.
2.2 Particle interacting with continuous mass distribution
We now generalize the above results to a particle interacting with a continuous mass distribution. Consider placing the point mass m at the origin, and it interacts with a mass $M$ distributed in space within a volume $V$, with a position-dependent density $\rho(\mathbf{r})$. In that case, the force on $m$ due to a small volume element $dV$ within the extended body is
$d\mathbf{F}=-Jm\rho(\mathbf{r})rdV.$ (8)
@@ 2.2 Particle interacting with continuous mass distribution
Notice, due to the linear nature of the force law, there is no concern of singularities in allowing the particle of interest to be located within the larger mass distribution. From (9), it is clear that the shape, size, and orientation of the mass distribution have no effect on the motion of test mass $m$. The distributed mass does not need to be a rigid body, and it may rotate in any arbitrary manner, as long as it is rotating about its center of mass, which is how unconstrained bodies rotate [3, 20, 21].
2.3 Potential energy
The potential energy $U$, of the system is also found analytically. The potential energy of a system of particles attached via the properly chosen springs is
@@ 2.3 Potential energy
where the additional factor of $\frac{1}{2}$ corrects for double counting. The case $i=k$ does not contribute to the sum, so it need not be removed. Expanding the product,
@@
The last term is the expression for the location of the center of mass, which is set to be zero. The summations over mass alone in the first two terms gives the total mass $M$, so the first two terms are identical. The resulting potential energy
@@
Alternatively, we calculate the work to move the $N$ particles from the origin to their positions $\mathbf{R}_{i}$ from the force law (1) directly, with the assumption that the center of mass remains at the origin. The work done in moving a single particle is
@@ 2.4 Repulsive Case
where $\Omega\equiv\sqrt{-JM}$. The energy analysis is the same as in Section 2.2, with $J\rightarrow-J$. Unless $\mathbf{C}=0$, the particle escapes to infinity, with exponentially increasing velocity. The potential energy of a particle is always negative
@@ 3 Applications
Numerical techniques explicitly evolve the system forward in time given the current state of positions and velocities. It is pedagogically useful to show that this numerical technique produces results that match the analytic solutions of Section 2. Animating the trajectories using VPython is also instructive, as it effectively illustrates the “non-interaction” of nearby particles. This also demonstrates the benefits of the Euler-Cromer method, which conserves energy in oscillatory motion, while the Euler method does not. Numerical techniques are also needed if the system is subjected to external forces. Examples of these applications are described in Appendix B.
@@
In the case of an isolated system, the analytic solutions (7) and (16) give the position and velocity of any particle at arbitrary time. In the case of a system inside a container, a continuous collision algorithm is used to handle any particles that collide with the container during a given time step. Any particle colliding with the wall affects the rest of the masses, because the change in velocity of that single particle affects the center of mass. The effects of the collisions on the velocity of the center of mass are accumulated while the check for collisions is made. The analytic solutions are updated to account for the change in center of mass velocity. The effects of a moving wall are also incorporated. This collision algorithm is elaborated in Appendix A.
3.1 Thermodynamics
Here the attractive and repulsive systems are examined in a 3D spherical container. A spherical container is necessary because a rectangular prism will not thermalize the motion of the particles, as the motion along each of the Cartesian coordinates is independent. A collection of 4,000 particles is created with spherically symmetric initial positions and velocities. The masses are chosen at random to be between 4 and 5 base units.
Figure 2: This figure shows relationships between thermodynamic quantities for the attractive case. Temperature (a) and pressure (b) both decrease as volume is increased. While pressure and volume have a clearly-defined relationship, it does not follow a power law.
In contrast to the ideal gas model, systems with long-range interactions do not have well-defined thermodynamic quantities that apply to the entire system [22, 23, 24]. Instead, we examine the pressure and temperature at the boundary of the container. The pressure is calculated by summing over the momentum of the particles that hit the wall in each time step,
@@ 3.1 Thermodynamics
where $p_{r}$ is the radial component of the particle’s momentum R, is the radius of the container, and $\Delta t$ is the computational time step. The temperature is determined by averaging the kinetic energy of the particles that hit the wall in a given time interval. The pressure and temperature are therefore only defined in terms of the interaction with the environment via the walls of the container. This is similar to the definition of thermodynamic parameters for bound gravitational systems [22, 23, 24]. In the attractive case, pressure and temperature are well defined only when the container is sufficiently small that the particles interact with the walls of the container.
We change the container radius quickly, allowing the particles to undergo free expansion into the increased volume. The system is then given time after each free expansion to come back to equilibrium. The thermalization time is particularly long for the repulsive case, as expanding the container changes the kinetic energy of the system notably. The values presented in the Figures 2 and 3 are the values at the end of the time segment used to equilibrate. All quantities are scaled by their values after an initial thermalization period. It is worthwhile to note that the only quantities being held fixed in these simulations is the number of particles and the total energy of the system.
@@
Figure 3: Relationships between thermodynamic quantities for the repulsive case. (a) Because the kinetic energy increases without bound, temperature increases with volume. (b) The repulsive force increases as the particles get farther apart, so pressure increases as the container’s volume increases.
@@
In examining Figures 2-3, it is evident that there is more scatter in the temperature values than in the pressure values. This is because the pressure depends on the radial component of the velocity of the particles that hit the wall, which will equilibrate due to collisions with the spherical container. The temperature also depends on the tangential component of the velocity, which is unaffected by contact with the wall.
@@
For the attractive case, the relationships among pressure, temperature, and volume share some similarity with those for the ideal gas model (see Figure 2). However, log-log plots reveal the relationships between pressure and volume and pressure and temperature do not follow power laws.
@@
For the repulsive case, an increase in the volume allows the potential energy to grow very large and negative, so that the kinetic energy of the particles impacting the wall of the container increases. As such, Figure 3 shows that the pressure and the temperature both increase with increasing volume, contrary to the behavior of attractive systems and ideal gases. The relationship between pressure and volume appears linear on visual inspection of Figure 3(b), but a log-log fit reveals that $P\sim R^{2.5}$ for this data.
4 Conclusions
@@
We present a pair-wise $N$-body interaction that is shown to be equivalent to $N$ one-body interactions with the center of mass of the system. In both attractive and repulsive case, the analytical trajectories are given by harmonic oscillator solutions. This is in stark contrast to other long-range interactions such Newtonian gravity, where only special case solutions exist for $N>2$.
@@
The mathematical form of the solutions, indicating the independent classical motion of each of the masses, implies a separable Hamiltonian. It has been known for some time that, with identical particles, the attractive case leads to a separable Hamiltonian. [13] The results presented here show that such a solution for the quantum equivalent [11, 12, 13, 14, 15] is also possible for N particles with arbitrary masses. The repulsive case is not bound, so there is not an equivalent quantum mechanical model system.
@@ Appendix A Collision handling algorithm for particles in a container
This algorithm uses event-driven collision detection and elastic collision handling to simulate interactions between particles and the walls of a container. This is a discrete collision detection algorithm [25, 26], as opposed to the more computationally expensive continuous collision detection algorithm [27].
@@
Using the updated center of mass velocity and the current positions and velocities of the particles, new parameters for each particle’s trajectories are calculated. The positions and velocities of all particles are then advanced to $t+\Delta t$, ensuring that no further collisions occur within this time step. The algorithm then repeats this process for subsequent time steps.
@@
As an example, the data provided involves $N=$ 40,000 particles over 500,000 time steps. This simulation runs on a standard laptop in less than 40 seconds. For scenarios involving fixed walls, or when expansion into a larger container is permitted, energy conservation is confirmed within computational precision.
Appendix B Suggested student projects

### v3

@@ Abstract
We present a pair-wise force law in a system of $N$ particles that produces analytic solutions for an arbitrary number of particles, masses, and initial conditions. Each pair of particles interacts via a force that is proportional to the product of their masses and their separation distance, with the force directed radially. We show that, despite the $N$-body interaction, each particle behaves as if it interacts only with the center of mass of the system. This effective one-body interaction behaves as Hooke’s Law with a common frequency for all particles, with the familiar analytic solutions for the trajectories. With these analytic solutions, it is possible to efficiently simulate a collection of these particles and incorporate other external forces. As an example, we simulate the particles within an adiabatically expanding container and calculate pressure and temperature in both the attractive and repulsive cases.
I Introduction
@@
The general $N$-body problem has long been a challenge in analytical mechanics. Newtonian gravity has been of particular interest, where the three-body problem admits analytical solutions only in special cases. Sundman1913 ; Moore1993 ; Chenciner2000 ; Li2018 ; Przybylska2023 ; uvakov2014 ; Delshams2018 ; Wild1980 ; Harmon2003 ; Khandekar1972 Even with numerical solutions, the chaotic nature of $N$-body problems means that arbitrary systems are unstable, presenting pedagogical challenges for students interested in simulating such a system.
@@
This paper presents an investigation of a pairwise force law with analytic solutions in an $N$-body system. Intriguingly, this force law is shown to be equivalent to a central force acting on each particle, and each individual particle exhibits motion independent of its immediate neighbors. Given their mathematical simplicity, these solutions can be conveniently integrated into undergraduate courses on advanced mechanics, within units on harmonic oscillators or gravitational interactions. Moreover, due to their computational efficiency, these systems can be accurately simulated with large $N$ values on personal computers, thus rendering them suitable for inclusion in computational projects or courses on numerical methods.
@@
The absence of short range interactions is convenient for quantum systems based on similar interactions, as the Hamiltonian is separable Calogero1969 ; Calogero1971 ; Lim1971 ; Vladimiroff1974 ; BialynickiBirula1985 ; Crandall1984 . However, this simplicity means that the classical system in isolation lacks the interesting orbital dynamics of Newtonian gravity. For this reason, further applications of this force law are most intriguing when paired with an external interaction. In this paper, we explore a thermodynamic application with the system in a container and suggest follow-up student projects.
@@
The paper is organized as follows. In Section II the force law is introduced and the equivalence of an $N$-body interaction to a one-body interaction with the center of mass is shown. In Section III, we present thermodynamic results for attractive and repulsive systems as a function of volume. Conclusions are provided in Section IV. The collision algorithm is presented in Appendix A, while ideas for possible research projects for advanced undergraduate students involving these two systems are discussed in Appendix B. Supplemental materials include a number of short python programs, including one for plotting the orbits of an N-particle system given a specified set of masses and initial conditions based on the analytic trajectories. A similar program that runs a numerical simulation of a similar system directly from the pair-wise forces is also included.
II Force Law and General Solution
@@
Bertrand’s theorem shows that the only central forces that guarantee closed orbits are the inverse square and linear force laws. bertrand1873 Inspired by this, we introduce a toy model of a linear gravity-like force
@@
II.1 Solutions for point particles
@@
where $M=\sum_{k}m_{k}$ is the total mass of the system and $\mathbf{R}_{\mathrm{cm}}$ is the position of the center of mass, which need not be stationary. Thus, the particle behaves as if it is governed by Hooke’s Law centered at $\mathbf{R}_{\mathrm{cm}}$, with effective spring constant $k_{i}=JMm_{i},$ as illustrated in Fig. 1.
@@
where $\omega=\sqrt{JM}$ for all particles and the constant vectors $\mathbf{A},\,\mathbf{B}$ are determined as usual from initial conditions. In the absence of external forces, the center of mass of the system moves at constant velocity, and we can use coordinates with $\mathbf{R}_{\mathrm{cm}}=0$ so that the particles move in closed elliptical orbits with a common frequency, consistent with Bertrand’s Theorem. bertrand1873 ; taylor2005classical ; goldstein2001classical ; grossman1996sheer
Thus, the $N$-body interaction has an analytical solution where the motion of each particle is effectively described as an interaction with the center of mass of the system. This differs from the general $N$-body problem, which cannot be solved analytically. Additionally, particles in this system remain bound, in contrast to the inverse-square law system, which can eject particles from the system.
II.2 Particle interacting with continuous mass distribution
We now generalize the above results to a particle interacting with a continuous mass distribution. Consider a point mass m at the origin that interacts with a mass $M$ distributed in space within a volume $V$ that has a position-dependent density $\rho(\mathbf{r})$. In that case, the force on $m$ due to a small volume element $dV$ within the extended body is
$d\mathbf{F}=-Jm\left(\rho(\mathbf{r)}dV\right)\mathbf{r}.$ (8)
@@ II.2 Particle interacting with continuous mass distribution
Notice, due to the linear nature of the force law, there is no concern of singularities in allowing the particle of interest to be located within the larger mass distribution. From (9), it is clear that the shape, size, and orientation of the mass distribution have no effect on the motion of test mass $m$. The distributed mass does not need to be a rigid body, and it may rotate in any arbitrary manner, as long as it is rotating about its center of mass, which is how unconstrained bodies rotate. fowles1999analytical ; Thornton_Stephen_T_2022-01-03 ; Hamill2022
II.3 Potential energy
The potential energy $U$ of the system is also found analytically. The potential energy of a system of particles attached via the properly chosen springs is
@@ II.3 Potential energy
where the additional factor of $\frac{1}{2}$ corrects for double counting. The case $i=k$ does not contribute to the sum, so it need not be removed. Expanding the product yields
@@
The last term is proportional to the squared magnitude of the location of the center of mass, which is set to be zero. The summations over mass alone in the first two terms gives the total mass $M$, so the first two terms are identical. The resulting potential energy
@@
Alternatively, we calculate the work to move the $N$ particles from the origin to their positions $\mathbf{r}_{k}$ from the force law (1) directly. The work done in moving a single particle is
@@ II.4 Repulsive Case
where $\Omega\equiv\sqrt{-JM}$. The energy analysis is the same as in Section II.2, with $J\rightarrow-J$. Unless $\mathbf{C}=0$, the particle escapes to infinity with exponentially increasing velocity. The potential energy of a particle is always negative
@@ III Applications
Numerical techniques explicitly evolve the system forward in time given the current state of positions and velocities. It is pedagogically useful to show that this numerical technique produces results that match the analytic solutions of Section II. Animating the trajectories using VPython is also instructive, as it effectively illustrates the “non-interaction” of nearby particles. This also demonstrates the benefits of the Euler-Cromer method, which conserves energy in oscillatory motion, while the Euler method does not. Numerical techniques are also needed if the system is subjected to external forces. Examples of these applications are described in Appendix B.
@@
In the case of an isolated system, the analytic solutions (7) and (16) give the position and velocity of any particle at arbitrary time. In the case of a system inside a container, a continuous collision algorithm is used to handle any particles that collide with the container during a given time step. Any particle colliding with the wall affects the rest of the masses, because the change in velocity of that single particle affects the center of mass. The effects of the collisions on the velocity of the center of mass are accumulated while the check for collisions is made. The analytic solutions are updated to account for the change in center of mass velocity. The effects of an adiabatically expanding container are also incorporated. This collision algorithm is elaborated in Appendix A.
III.1 Thermodynamic relationships
Figure 2: This figure shows relationships between thermodynamic quantities for the attractive case. Temperature (a) and pressure (b) both decrease as volume is increased. Plot (c) shows the resulting positive correlation between pressure and temperature. From visual inspection, one might expect a power-law relationship among these quantities, but analysis of log-log plots indicates no such behavior.
Here the attractive and repulsive systems are examined in a 3D spherical container. A spherical container is necessary because a rectangular prism will not thermalize the motion of the particles, as the motion along each of the Cartesian coordinates is independent. A collection of 4,000 particles is created with randomized initial positions and velocities within the initial spherical container. The masses are chosen at random to be between 4 and 5 base units.
In contrast to the ideal gas model, systems with long-range interactions do not have well-defined thermodynamic quantities that apply to the entire system. BerberanSantos1997 ; Lente2020 ; MenesesJoo2022 Instead, we examine the pressure and temperature at the boundary of the container. The pressure is calculated by summing the change in momentum of the particles that hit the wall in each time step,
@@ III.1 Thermodynamic relationships
where $p_{r}$ is the radial component of the particle’s momentum R, is the radius of the container, and $\Delta t$ is the computational time step. The temperature is determined by averaging the kinetic energy of the particles that hit the wall in a given time interval. The pressure and temperature are therefore only defined in terms of the interaction with the environment via the walls of the container. This is similar to the definition of thermodynamic parameters for bound gravitational systems. BerberanSantos1997 ; Lente2020 ; MenesesJoo2022 In the attractive case, pressure and temperature are well defined only when the container is sufficiently small that the particles interact with the walls of the container.
The only quantity held fixed in these simulations is the number of particles. The discontinuous motion of the container and the elastic collisions ensures no work is done on the system, so total energy remains constant.
@@
Figure 3: Relationships between thermodynamic quantities for the repulsive case. (a) Because the kinetic energy increases without bound, temperature increases with volume. (b) Pressure is inversely proportional to the square root of volume. (c) In contrast to the attractive case, pressure and temperature are negatively correlated. The relationship between temperature and the other quantities cannot be described by a power law.
@@
In examining Figs. 2-3, it is evident that there is more scatter in the temperature values than in the pressure values. This is because the pressure depends on the radial component of the velocity of the particles that hit the wall, which will equilibrate due to collisions with the spherical container. The temperature also depends on the tangential component of the velocity, which is unaffected by contact with the wall.
@@
For the attractive case, the relationships among pressure, temperature, and volume share some similarity with those for the ideal gas model (see Fig. 2). However, log-log plots indicate the relationships do not follow power laws.
@@
For the repulsive case, an increase in the volume allows the potential energy to grow very large and negative, so that the kinetic energy of the particles impacting the wall of the container increases. As such, Fig. 3 shows that the temperature increases with increasing volume, contrary to the behavior of attractive systems and ideal gases. However, pressure decreases as a result of the increasing surface area, and we find $P\sim V^{1/2}$ from a log-log fit of the data. The relationship between pressure and volume appears linear on visual inspection of Fig. 3(b), but a log-log fit reveals that the relationships involving temperature do not follow power laws.
III.2 Thermodynamic equilibration
In the simulation, the container is expanded adiabatically in a series of discontinuous jumps, and the particles are allowed some time for thermodynamic equilibration after each expansion. The container radius is increased discontinuously with a time interval $\tau$ between increases, allowing the particles to undergo free expansion into the increased volume. Thermalization is allowed for a duration $0.8\tau$, and the thermodynamic quantities are averaged over the final $0.2\tau$ before the subsequent expansion. The thermalization time is particularly long for the repulsive case, as expanding the container changes the kinetic energy of the system. The values presented in Figs. 2 and 3 are the values at the end of the time segment used to equilibrate.
In Fig. 4, we show a simulation of the attractive system in a fixed container over a short period of time $\omega t=10/\pi$, plotting the values at each time step, with no averaging. The particles start out in a random configuration, but after interactions with the container walls, the pressure and temperature settle down toward equilibrium values.
Figure 4: Thermodynamic equilibration over a short time period in a fixed volume is illustrated by the pressure (left) and temperature (right). The values are normalized by the average of the data for $\omega t>6$.
We compare the effects of equilibration time in Fig. 5. In the short equilibration simulation, the equilibration time is $\sim 12.7$ times the oscillation period of the particles. The long equilibration is 100 times longer. When there is less equilibration time, the particles may not settle to the equilibrium state, and there is more noise in the data, although the same overall trend is present.
Figure 5: A comparison of results with different thermal equilibration times. The solid blue lines represent the same results shown in Fig. 2. The orange dashed lines represent simulations where the container is expanded in the same small increments, but with less time for the particles to reach thermal equilibrium. The overall trend is still visible, but the relationships are less clear.
IV Conclusions
@@
In this paper, we showed that a pair-wise $N$-body interaction is equivalent to $N$ one-body interactions with the center of mass of the system. In the attractive and repulsive case, the analytical trajectories are given by harmonic oscillator and exponential solutions, respectively. This is in stark contrast to other long-range interactions such Newtonian gravity, where only special case solutions exist for $N>2$.
@@
The mathematical form of the solutions, indicating the independent classical motion of each of the masses, implies a separable Hamiltonian, which has been investigated in quantum systems of identical particles. Calogero1969 ; Calogero1971 ; BialynickiBirula1985 ; Lim1971 ; Vladimiroff1974 The results presented here show that such a solution is possible in a quantum equivalent with differing masses.
The authors are actively working to extend the well-known systems of simple harmonic oscillator pair-wise interacting quantum systems of identical masses Calogero1971 ; Lim1971 ; Vladimiroff1974 ; Kocik2013 ; Chen2018 ; Sowiski2019 ; ZauskaKotur2000 , to cases where the masses of the particles are not necessarily identical. In particular, a system with mixed masses seems very promising as an educational tool as a relatively simple model of the strong force and nuclear physics.
@@ Appendix A Collision handling algorithm for particles in a container
This algorithm uses event-driven collision detection and elastic collision handling to simulate interactions between particles and the walls of a container. This is a discrete collision detection algorithm, GarciaAlonso1994 ; Uchiki1983 as opposed to the more computationally expensive continuous collision detection algorithm. Redon2002
@@
Using the updated center of mass velocity and the current positions and velocities of the particles, new parameters for each particle’s trajectories are calculated. The positions and velocities of all particles are then advanced to $t+\Delta t$, ensuring that no further collisions occur within this time step. The algorithm then repeats this process for subsequent time steps. This collision-handling algorithm has an $\mathcal{O}(\Delta t)$ error in both position and velocity.
@@
As an example, the data provided in Figs. 2 and 3 involves $N=$ 40,000 particles over 500,000 time steps. This simulation runs on a standard laptop in less than 40 seconds. For scenarios involving fixed walls, or when expansion into a larger container is permitted, energy conservation is confirmed within computational precision.
Appendix B Student Projects